# Fully Dynamic Euclidean $k$-Means

### v2

@@ 3.1 Robust Centers
Definition 3.1 (Robust Center, Definition 3.2 in the arxiv version of [BCF25a]).
Assume $X\subseteq[\Delta]^{d}$. Let $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points in $[\Delta]^{d}$, and let $B_{i}=\textsc{Ball}_{X}(x_{i},10^{i})$ for each $i\in[0,t]$. We refer to $(x_{0},x_{1},\ldots,x_{t})$ as a $t$-robust sequence w.r.t. $X$ iff for every $i\in[1,t]$:
@@
Definition 3.2 (Robust Solution, Definition 3.5 in the arxiv version of [BCF25a]).
@@
$u\ \text{is}\ t\text{-robust w.r.t.\ $X$, where}\ t\ \text{is the smallest integer satisfying}\ 10^{t}\geq\mathrm{dist}(u,S-u)/200.$ (1)
@@ 3.2 Description of The Algorithm
Barriers to Implementation in Euclidean Spaces. To implement this algorithmic framework in general metric spaces, [BCF25a] give various different subroutines for each of these steps efficiently. In particular, they design algorithms for solving the ‘restricted’ and ‘augmented’ $k$-means problems (see Section 4.4) and implement the Robustify subroutine (see Section 5.3). While these subroutines are fast enough for general metric spaces (where we can only hope to get an update time of $\tilde{O}(k)$), there are too slow for our setting (where we can obtain an update time of $\tilde{O}(k^{\epsilon})$). For example, the time taken to implement Step 1 and determine the length of an epoch in [BCF25a] is $\tilde{O}(k\cdot(\ell+1))$. Since this can be amortized over an epoch of length $\ell$, this leads to $\tilde{O}(k)$ amortized update time. In the Euclidean setting, however, we would need to perform this task in $\tilde{O}(k^{\epsilon}\cdot(\ell+1))$ time to get our desired bounds. Thus, we must develop new geometric data structures that allow us to implement versions of these subroutines much more efficiently in Euclidean spaces. In Section 4, we present the new geometric data structures that we introduce for our purpose.
@@ 4 Data Structures in High-Dimensional Euclidean Spaces
In this section, we present new fundamental data structures required to implement the framework of [BCF25a] in high-dimensional Euclidean spaces. There are some parameters we use throughout this section, whose final values in our algorithm are being set as follows.
@@
Our data structures rely on a fundamental geometric tool called consistent hashing [JLN+05, CJK+22, Fil24], specifically $(\Gamma,\Lambda)$-hashes as in Definition 4.1. We provide the first time-efficient consistent hashing with competitive parameters, strengthening previous efficient constructions that only achieve a relaxed guarantee (see Section 4.1). Then in Sections 4.2, 4.3 and 4.4, we leverage our new consistent hashing scheme to design several dynamic data structures against adaptive adversaries, especially to aid the two key sub-tasks called restricted $k$-means and augmented $k$-means problems.
Although our final application is clustering, the structures we devise may be useful for more general tasks. Specifically, we devise a reduction-style approximate range query structure, that turns any structure whose query concerns only global dataset, into one that supports range query, and a typical application is to return the coreset [HM04] with respect to a ball range given as query. We also devise a structure that approximately maintains the approximate nearest-neighbor assignment between a dynamic dataset and a dynamic center set, which may be used to approximately draw $D^{2}$-samples [AV07].
In this section, when we say “with high probability” (w.h.p.), we mean the algorithm satisfies all stated guarantee simultaneously with probability $1-1/\mathrm{poly}(\Delta^{d})$. We assume the access to a $(\Gamma,\Lambda)$-hash family in Sections 4.2, 4.3 and 4.4, where we use the $\Gamma$ and $\Lambda$ parameters without further quantification.
@@ 4.1 Efficient Consistent Hashing
We give in Definition 4.1 the notion of consistent hashing, which was introduced in [CJK+22] and was also studied under a closely related notion called sparse partitions [JLN+05, Fil24]. Roughly speaking, a consistent hashing is a data-oblivious map on $[\Delta]^{d}$ that partitions the space $[\Delta]^{d}$ into hash buckets, such that each bucket has some bounded diameter $\rho$, and every small ball of diameter $\frac{\rho}{\Gamma}$ can intersect only $\Lambda$ parts.
@@
A map $\varphi$ on $[\Delta]^{d}$ is a $(\Gamma,\Lambda,\rho)$-hash if
@@
We also need the hash function to evaluate efficiently. In particular, given a point $x\in[\Delta]^{d}$, we need to evaluate $\varphi(x)$ and $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ (allowing approximation) in $O(\Lambda)$ time (noticing that $|\varphi(\textsc{Ball}(x,\rho/\Gamma))|\leq O(\Lambda)$ by the consistency guarantee), as follows.
@@
We devise efficient consistent hashing with parameter $\Lambda=2^{O(d/\Gamma^{2/3})}\cdot\mathrm{poly}(d)$. This tradeoff is close to the optimal existential bound of $\Lambda=2^{\Omega(d/\Gamma)}$ from [Fil24] (which runs in $\exp(d)$ time). Previously, efficient construction of consistent hashing is only known from [FJL+25], but their consistency guarantee is in expectation which is too weak to satisfy Definition 4.1. While such a weaker guarantee may be useful in many applications, it is nonetheless insufficient for our application especially for devising structures robust to adaptive adversary.
@@
For any $\Gamma\geq 2\sqrt{2\pi}$, $\rho>0$ and $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, there exists a random hash $\varphi$ on $[\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-1/\mathrm{poly}(\Delta^{d})$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
@@
Construction. We sketch the proof of Lemma 4.3. Our starting point is the above-mentioned weaker construction from [FJL+25] which only achieves an expected consistency guarantee. We aim to boost this expected guarantee to a high-probability one with only $\tilde{O}(1)$ loss on the consistency bound. Specifically, let $\tilde{\varphi}$ be a random hashing on $\mathbb{R}^{d}$, the expected consistency guarantee of $\tilde{\varphi}$ is
@@
$\forall x\in[\Delta]^{d},\quad\mathbb{E}\big{[}|\tilde{\varphi}(\textsc{Ball}(x,{\rho}/{\Gamma}))|\big{]}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \Lambda.$
@@
We describe our construction. We first draw $m=O(d\log\Delta)$ independent weaker hashing $\{\tilde{\varphi}_{i}\}_{i\in[m]}$ [FJL+25] with expected consistency guarantee. Then, with high probability, for all points $x\in[\Delta]^{d}$, there exists $i\in[m]$ such that the strict consistency condition $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))|\leq 2\Lambda$ holds. We pick for each point $x$ the smallest such $i\in[m]$ and call this the color of $x$, denoted as $c_{x}$. Our hash $\varphi$ is defined as $\varphi(x):=(c_{x},\tilde{\varphi}_{c_{x}}(x))$, $\forall x\in[\Delta]^{d}$.
@@
Consistency. The diameter bound $\mathrm{diam}(\varphi^{-1}(z))\leq\rho$ carries over from $\tilde{\varphi}$’s. We next show the consistency $|\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq O(m\Gamma)$ for $x\in[\Delta]^{d}$. Let $Y:=\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))$, and we need $|Y|\leq O(\Lambda)$. For $i\in[m]$ let $Y_{i}:=\{(c_{x},\tilde{\varphi}_{c_{x}}(x))\in Y:c_{x}=i\}$ be the subset of hash value coming from color $i$. Then $|Y|\leq\sum_{i}|Y_{i}|$. We show that $|Y_{i}|\leq O(\Gamma)$, and this leads to $|Y|\leq O(m\Gamma)$ as desired.
@@
Fix some $i$. By our construction, an immediate bound is $|Y_{i}|\leq|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|$, so we are done if $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Gamma$. However, it could be that $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|>2\Gamma$. We show that if this happens, then no point in $\textsc{Ball}(x,\rho/(2\Gamma))$ can have color $i$, implying $|Y_{i}|=0$. This follows from the contrapositive: suppose some $y\in\textsc{Ball}(x,\rho/(2\Gamma))$ has color $c_{y}=i$. Then $\textsc{Ball}(x,\rho/(2\Gamma))\subseteq\textsc{Ball}(y,\rho/\Gamma)$, so the consistency guarantee at $y$ implies $|\phi_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Lambda$.
@@
Efficiency. Our new construction makes the hashing more complicated to evaluate, particularly to determine the color $c_{x}$ for $x\in[\Delta]^{d}$. By the definition of $c_{x}$, this reduces to checking whether $N_{i}:=|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))|\leq 2\Lambda$ for each color $i$. To do this efficiently, we leverage the special structure of $\tilde{\varphi}_{i}$ in [FJL+25]: each $\tilde{\varphi}_{i}$ is defined using a randomly shifted grid over $[\Delta]^{d}$, where each point is hashed based on the grid cell that it belongs to. Under this grid structure, $\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))$ corresponds to a connected set of grid cells, where each cell is adjacent to $O(d)$ others (i.e., differs in only one coordinate). We can explore these cells using a depth-first search, and once the depth-first search takes more than $\tilde{O}(\Lambda)$ time, we can terminate the search and conclude that $N_{i}>2\Lambda$, and no point around $x$ has color $i$. This entire process takes total time $N_{i}\cdot\mathrm{poly}(d)$. The computation of $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ follows similar ideas.
@@ 4.2 Approximate Range Query Structure
An immediate application of consistent hashing yields a range query structure. This works as a general reduction, that turns a data structure $\mathcal{A}$ which answers query regarding only the entire dataset, to one that can answer queries on a metric ball $\textsc{Ball}(x,r)$. Specifically, we consider algorithm $\mathcal{A}$ that maintains dynamic dataset $X\subseteq[\Delta]^{d}$, and that it reports some element $e_{X}$ on some domain $U$. For example, elements in $U$ may be real numbers, or a subset of points. We say an element $e_{T}\in U$ admissible to $\mathcal{A}$ with respect to some dataset $T$, if it satisfies the specification of $\mathcal{A}$, such as having some approximation ratio $\beta$ on $T$.
@@
Suppose an algorithm $\mathcal{A}$ maintains dynamic $X\subseteq[\Delta]^{d}$ and it reports $e_{X}\in U$ as query. There exists a dynamic algorithm that takes $r>0$ and dynamic $X\subseteq[\Delta]^{d}$ as input, and for every query $x\in[\Delta]^{d}$ reports $\{e_{1},\ldots,e_{s}\}$ for $s\leq O(\Lambda)$ such that there exists an $s$-partition $\{T_{1},\ldots,T_{s}\}$ of $\textsc{Ball}^{\prime}(x,r)\cap X$ that satisfies $\textsc{Ball}(x,r)\subseteq\textsc{Ball}^{\prime}(x,r)\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)$, and $e_{i}$ is admissible with respect to $T_{i}$ ($\forall i\in[s]$). The update and query run in $\tilde{O}(\Lambda)$ times that of $\mathcal{A}$, respectively. This algorithm works against adaptive adversary if $\mathcal{A}$ does, and succeeds w.h.p.
Application: Range Query Coreset. This reduction is useful when the query output of $\mathcal{A}$ can be merged, i.e., $e_{1},\ldots,e_{s}$ can be used to derive $e_{\textsc{Ball}^{\prime}(x,r)}$ for dataset $\textsc{Ball}^{\prime}(x,r)\cap X$. In particular, we devise a structure that reports a coreset of size $O(k)$ for $k$-means with respect to $T$ for every query $x\in[\Delta]^{d}$. Here, a coreset with respect to some $T\subseteq[\Delta]^{d}$ is a (weighted) subset $D$ such that $\forall S\subseteq[\Delta]^{d},|S|\leq k$ one has $\textsc{cost}(D,S)=\Theta(1)\cdot\textsc{cost}(T,S)$ [HM04]. We can implement this coreset with range query by plugging in Lemma 4.4 a known dynamic coreset algorithm [HK20, lTHS24] as $\mathcal{A}$. Then we can obtain a coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ via a standard merge-and-reduce framework [HM04]. In our application, we need this coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ with $k=1$, so the size becomes $O(1)$, which enables us to compute approximate solutions and cost estimations over the approximate ball $\textsc{Ball}^{\prime}(x,r)$ very efficiently, which is stated in the following lemma.
@@
There is an algorithm that maintains dynamic dataset $X\subseteq[\Delta]^{d}$, such that given a point $x\in[\Delta]^{d}$ and a radius $r\geq 0$, we can compute in time $\tilde{O}(\Lambda)$ a real number $b_{x}\geq 0$, a center $c^{\star}\in[\Delta]^{d}$, and two estimations $\widehat{\textsc{cost}}_{c^{\star}}$ and $\widehat{\textsc{cost}}_{x}$ such that there exists some point subset $B_{x}\subseteq X$ with $\textsc{Ball}(x,r)\subseteq B_{x}\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)$ satisfying the following.
@@
The algorithm has amortized update time $\tilde{O}(\Lambda)$, works against an adaptive adversary, and succeeds w.h.p.
Application: ANN. In addition, Lemma 4.4 can also be applied to yield an $O(\Gamma)$-approximate nearest neighbor (ANN) structure with update and query time $\tilde{O}(\Lambda)$, against adaptive adversary. Similar guarantee of ANN has also been obtained in e.g., [BDF+24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing. We summarize this guarantee as the following lemma which is needed in our main algorithm.
@@
The data structure has amortized update time $\tilde{O}(\Lambda)$, works against an adaptive adversary, and succeeds w.h.p.
Proof Sketch of Lemma 4.4. The algorithm starts with imposing a $(\Gamma,\Lambda,\rho:=\Gamma r)$-consistent hashing $\varphi$ (Lemma 4.3), and run an independent instance of $\mathcal{A}$ for each bucket. To answer the query, one compute $\varphi(\textsc{Ball}(x,r))$, and return the query result of $\mathcal{A}$ for each bucket $z\in\varphi(\textsc{Ball}(x,r))$. The analysis is straightforward and follows easily from Definition 4.2. Previously, an MPC version of Lemma 4.4 was introduced in [CGJ+24], but their algorithm is not efficient due to the lack of efficient consistent hashing which we devise (Lemma 4.3).

### v3

@@ 4.1 Robust Centers
Definition 4.1 (Robust Center, Definition 3.2 in the arxiv version of [BCF24]).
Assume $X\subseteq[\Delta]^{d}$. Let $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points in $\Delta]^{d}$, and let $B_{i}=\textsc{Ball}_{X}(x_{i},10^{i})$ for each $i\in[0,t]$. We refer to $(x_{0},x_{1},\ldots,x_{t})$ as a $t$-robust sequence w.r.t. $X$ iff for every $i\in[1,t]$:
@@
Definition 4.2 (Robust Solution, Definition 3.5 in the arxiv version of [BCF24]).
@@
[…]
@@ 4.2 Description of The Algorithm
At the start of an epoch, the set $S$ is robust and $\textsc{cost}(X,S)\leq O(1)\cdot\mathrm{OPT}_{k}(X)$.
@@
Barriers to Implementation in Euclidean Spaces. To implement this algorithmic framework in general metric spaces, [BCF24] gives various different subroutines for each of these steps efficiently. In particular, they design algorithms for solving the ‘restricted’ and ‘augmented’ $k$-means problems (see Section 5.4) and implement the Robustify subroutine (see Section 6.3). While these subroutines are fast enough for general metric spaces (where we can only hope to get an update time of $\tilde{O}(k)$), they are too slow for our setting (where we can obtain an update time of $\tilde{O}(k^{\epsilon})$). For example, the time taken to implement Step 1 and determine the length of an epoch in [BCF24] is $\tilde{O}(k\cdot(\ell+1))$. Since this can be amortized over an epoch of length $\ell$, this leads to $\tilde{O}(k)$ amortized update time. In the Euclidean setting, however, we would need to perform this task in $\tilde{O}(k^{\epsilon}\cdot(\ell+1))$ time to get our desired bounds. Thus, we must develop new geometric data structures that allow us to implement versions of these subroutines much more efficiently in Euclidean spaces. In Section 5, we present the new geometric data structures that we introduce for our purpose.
@@ 5 Data Structures in High-Dimensional Euclidean Spaces
In this section, we present new fundamental data structures required to implement the framework of [BCF24] in high-dimensional Euclidean spaces. There are some parameters we use throughout this section, whose final values in our algorithm are being set as follows.
@@
Our data structures rely on a fundamental geometric tool called consistent hashing [JiaLNRS05, CJKVY22, Filtser24], specifically $(\Gamma,\Lambda)$-hashes as in Definition 5.1. We provide the first time-efficient consistent hashing with competitive parameters, strengthening previous efficient constructions that only achieve a relaxed guarantee (see Section 5.1). Then in Sections 5.2, 5.3 and 5.4, we leverage our new consistent hashing scheme to design several dynamic data structures against adaptive adversaries, especially to aid the two key sub-tasks called restricted $k$-means and augmented $k$-means problems.
Although our final application is clustering, the structures we devise may be useful for more general tasks. Specifically, we devise a reduction-style approximate range query structure, that turns any structure whose query concerns only global dataset, into one that supports range query, and a typical application is to return the coreset [DBLP:conf/stoc/Har-PeledM04] with respect to a ball range given as query. We also devise a structure that approximately maintains the approximate nearest-neighbor assignment between a dynamic dataset and a dynamic center set, which may be used to approximately draw $D^{2}$-samples [ArthurV07].
In this section, when we say “with high probability” (w.h.p.), we mean the algorithm satisfies all stated guarantee simultaneously with probability $1-1/\mathrm{poly}(\Delta^{d})$. We assume the access to a $(\Gamma,\Lambda)$-hash family in Sections 5.2, 5.3 and 5.4, where we use the $\Gamma$ and $\Lambda$ parameters without further quantification.
@@ 5.1 Efficient Consistent Hashing
We give in Definition 5.1 the notion of consistent hashing, which was introduced in [CJKVY22] and was also studied under a closely related notion called sparse partitions [JiaLNRS05, Filtser24]. Roughly speaking, a consistent hashing is a data-oblivious map on $\Delta]^{d}$ that partitions the space $\Delta]^{d}$ into hash buckets, such that each bucket has some bounded diameter $\rho$, and every small ball of diameter $\frac{\rho}{\Gamma}$ can intersect only $\Lambda$ parts.
@@
A map $\varphi$ on $\Delta]^{d}$ is a $(\Gamma,\Lambda,\rho)$-hash if
@@
We also need the hash function to evaluate efficiently. In particular, given a point $x\in[\Delta]^{d}$, we need to evaluate $\varphi(x)$ and $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ (allowing approximation) in $\tilde{O}(\Lambda)$ time (noticing that $|\varphi(\textsc{Ball}(x,\rho/\Gamma))|\leq\Lambda$ by the consistency guarantee), as follows.
@@
We devise efficient consistent hashing with parameter $\Lambda=2^{O(d/\Gamma^{2/3})}\cdot\mathrm{poly}(d)$. This tradeoff is close to the optimal existential bound of $\Lambda=2^{\Omega(d/\Gamma)}$ from [Filtser24] (which runs in $\exp(d)$ time). Previously, efficient construction of consistent hashing is only known from [random-shift], but their consistency guarantee is in expectation which is too weak to satisfy Definition 5.1. While such a weaker guarantee may be useful in many applications, it is nonetheless insufficient for our application especially for devising structures robust to adaptive adversary.
@@
For any $\Gamma\geq 2\sqrt{2\pi}$, $\rho>0$ and $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, there exists a random hash $\varphi$ on $\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-1/\mathrm{poly}(\Delta^{d})$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
@@
Construction. We sketch the proof of Lemma 5.3. Our starting point is the above-mentioned weaker construction from [random-shift] which only achieves an expected consistency guarantee. We aim to boost this expected guarantee to a high-probability one with only $\tilde{O}(1)$ loss on the consistency bound. Specifically, let $\tilde{\varphi}$ be a random hashing on $\mathbb{R}^{d}$, the expected consistency guarantee of $\tilde{\varphi}$ is
@@
$\forall x\in[\Delta]^{d},\quad\mathbb{E}\big[|\tilde{\varphi}(\textsc{Ball}(x,{\rho}/{\Gamma}))|\big]\penalty 10000\ \leq\penalty 10000\ \Lambda.$
@@
We describe our construction. We first draw $m=O(d\log\Delta)$ independent weaker hashing $\{\tilde{\varphi}_{i}\}_{i\in[m]}$ [random-shift] with expected consistency guarantee. Then, with high probability, for all points $x\in[\Delta]^{d}$, there exists $i\in[m]$ such that the strict consistency condition $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))|\leq 2\Lambda$ holds. We pick for each point $x$ the smallest such $i\in[m]$ and call this the color of $x$, denoted as $c_{x}$. Our hash $\varphi$ is defined as $\varphi(x):=(c_{x},\tilde{\varphi}_{c_{x}}(x))$, $\forall x\in[\Delta]^{d}$.
@@
Consistency. The diameter bound $\mathrm{diam}(\varphi^{-1}(z))\leq\rho$ carries over from $\tilde{\varphi}_{i}$’s. We next show the consistency $|\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq O(m\Lambda)$ for $x\in[\Delta]^{d}$. Let $Y:=\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))$, and we need $|Y|\leq O(m\Lambda)$. For $i\in[m]$, let $Y_{i}:=\{(c_{x},\tilde{\varphi}_{c_{x}}(x))\in Y:c_{x}=i\}$ be the subset of hash values coming from color $i$. Then $|Y|\leq\sum_{i}|Y_{i}|$. We show that $|Y_{i}|\leq O(\Lambda)$, and this leads to $|Y|\leq O(m\Lambda)$ as desired.
@@
Fix some $i$. By our construction, an immediate bound is $|Y_{i}|\leq|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|$, so we are done if $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Lambda$. However, it could be that $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|>2\Lambda$. We show that if this happens, then no point in $\textsc{Ball}(x,\rho/(2\Gamma))$ can have color $i$, implying $|Y_{i}|=0$. This follows from the contrapositive: suppose some $y\in\textsc{Ball}(x,\rho/(2\Gamma))$ has color $c_{y}=i$. Then $\textsc{Ball}(x,\rho/(2\Gamma))\subseteq\textsc{Ball}(y,\rho/\Gamma)$, so the consistency guarantee at $y$ implies $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Lambda$.
@@
Efficiency. Our new construction makes the hashing more complicated to evaluate, particularly to determine the color $c_{x}$ for $x\in[\Delta]^{d}$. By the definition of $c_{x}$, this reduces to checking whether $N_{i}:=|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))|\leq 2\Lambda$ for each color $i$. To do this efficiently, we leverage the special structure of $\tilde{\varphi}_{i}$ in [random-shift]: each $\tilde{\varphi}_{i}$ is defined using a randomly shifted grid over $\Delta]^{d}$, where each point is hashed based on the grid cell that it belongs to. Under this grid structure, $\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))$ corresponds to a connected set of grid cells, where each cell is adjacent to $O(d)$ others (i.e., differs in only one coordinate). We can explore these cells using a depth-first search, and once the depth-first search takes more than $\tilde{O}(\Lambda)$ time, we can terminate the search and conclude that $N_{i}>2\Lambda$, and no point around $x$ has color $i$. This entire process takes total time $\tilde{O}(\Lambda)\cdot\mathrm{poly}(d)$. The computation of $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ follows from similar ideas.
@@ 5.2 Approximate Range Query Structure
An immediate application of consistent hashing yields a range query structure. This works as a general reduction, that turns a data structure $\mathcal{A}$ which answers query regarding only the entire dataset, to one that can answer queries on a metric ball $\textsc{Ball}(x,r)$. To model this, suppose $U$ is the domain for the output, then for each input dataset $T$, we identify a subset $U_{T}\subseteq U$ in the output domain as admissible output to the dataset $T$. For example, elements in $U$ may be real numbers, or a subset of points, and admissible output $U_{T}$ may consist of the solutions that have certain approximation ratio to the input $T$.
Under this modeling, we assume that the algorithm $\mathcal{A}$ generates admissible output for the entire input dataset, and our structure, stated as follows in Lemma 5.4, uses this algorithm in a black-box way to generate admissible outputs for a ball range query $\textsc{Ball}(x,r)$. Specifically, we would use consistent hashing technique to define some approximate $\textsc{Ball}^{\prime}(x,r)$ that is sandwiched between $\textsc{Ball}(x,r)$ (which is the desired ball) and $\textsc{Ball}(x,O(\Gamma)r)$ (which is the ball with slightly enlarged radius). By the consistent hashing guarantee, the approximate ball $\textsc{Ball}^{\prime}(x,r)$ may be represented as a disjoint union of $s:=O(\Lambda)$ buckets/subsets $T_{1},\ldots,T_{s}$. We then run $\mathcal{A}$ on $T_{i}$’s to generate admissible output $e_{1},\ldots,e_{s}$. We note that the following Lemma 5.4 stops at generating these admissible $e_{i}$’s, but to eventually obtain an admissible output for the (approximate) query $\textsc{Ball}^{\prime}(x,r)$, we still need a downstream algorithm that aggregates $e_{1},\ldots,e_{s}$. We choose not to include this step in Lemma 5.4 since this step may vary significantly for different applications, and we will discuss this aggregation step in concrete applications.
@@
Suppose we have an algorithm $\mathcal{A}$ on a dynamic dataset $X\subseteq[\Delta]^{d}$ and it reports some $e_{X}\in U$ that is admissible to $X$. Then there exists a dynamic algorithm that takes $r>0$ and dynamic $X\subseteq[\Delta]^{d}$ as input, and for each query point $x\in[\Delta]^{d}$ reports $e_{1},\ldots,e_{s}\in U$ that are admissible to some point sets $T_{1},\ldots,T_{s}$ respectively, where $s\leq O(\Lambda)$, such that $\{T_{1},\ldots,T_{s}\}$ is a an $s$-partition of some approximate ball $\textsc{Ball}^{\prime}(x,r)\cap X$ that satisfies $\textsc{Ball}(x,r)\subseteq\textsc{Ball}^{\prime}(x,r)\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)$. The update and query run in $\tilde{O}(\Lambda)$ times that of $\mathcal{A}$, respectively. This algorithm works against adaptive adversary if $\mathcal{A}$ does, and succeeds w.h.p.
Application: Range Query Coreset. This reduction is useful when the query output of $\mathcal{A}$ can be merged, i.e., $e_{1},\ldots,e_{s}$ can be used to derive $e_{\textsc{Ball}^{\prime}(x,r)}$ for dataset $\textsc{Ball}^{\prime}(x,r)\cap X$. In particular, we devise a structure that reports a coreset of size $O(k)$ for $k$-means with respect to $\textsc{Ball}^{\prime}(x,r)\cap X$ for every query $x\in[\Delta]^{d}$. Here, a coreset with respect to some $T\subseteq[\Delta]^{d}$ is a (weighted) subset $D$ such that $\forall S\subseteq[\Delta]^{d},|S|\leq k$ one has $\textsc{cost}(D,S)=\Theta(1)\cdot\textsc{cost}(T,S)$ [DBLP:conf/stoc/Har-PeledM04]. We can implement this coreset with range query by plugging in Lemma 5.4 a known dynamic coreset algorithm [HenzingerK20, DBLP:conf/esa/TourHS24] as $\mathcal{A}$. Then we can obtain a coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ via a standard merge-and-reduce framework [DBLP:conf/stoc/Har-PeledM04]. In our application, we need this coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ with $k=1$, so the size becomes $O(1)$, which enables us to compute approximate solutions and cost estimations over the approximate ball $\textsc{Ball}^{\prime}(x,r)$ very efficiently, which is stated in the following lemma.
@@
There is an algorithm on a dynamic dataset $X\subseteq[\Delta]^{d}$, such that given a point $x\in[\Delta]^{d}$ and a radius $r\geq 0$, it can compute in time $\tilde{O}(\Lambda)$ a real number $b_{x}\geq 0$, a center $c^{\star}\in[\Delta]^{d}$, and two estimations $\widehat{\textsc{cost}}_{c^{\star}}$ and $\widehat{\textsc{cost}}_{x}$ such that there exists some point subset $B_{x}\subseteq X$ with $\textsc{Ball}(x,r)\cap X\subseteq B_{x}\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)\cap X$ satisfying the following.
@@
The algorithm has amortized update time $\tilde{O}(\Lambda)$ (w.r.t. updates on $X$), works against an adaptive adversary, and succeeds w.h.p.
Application: ANN. In addition, Lemma 5.4 can also be applied to yield an $O(\Gamma)$-approximate nearest neighbor (ANN) structure with update and query time $\tilde{O}(\Lambda)$, against adaptive adversary. Similar guarantee of ANN has also been obtained in e.g., [DBLP:conf/nips/0001DFGHJL24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing. By plugging this ANN data structure in the algorithm of [BhattacharyaGJQ24], we obtain a new data structure that maintains, for each point, its distance to the approximate nearest neighbor. We summarize this result in the following lemma, which is needed in our main algorithm.
@@
The data structure has amortized update time $\tilde{O}(\Lambda)$ (w.r.t. updates on $S$), works against an adaptive adversary, and succeeds w.h.p.
Proof Sketch of Lemma 5.4. The algorithm starts with imposing a $(\Gamma,\Lambda,\rho:=\Gamma r)$-consistent hashing $\varphi$ (Lemma 5.3), and run an independent instance of $\mathcal{A}$ for each bucket. To answer the query, one compute $\varphi(\textsc{Ball}(x,r))$, and return the query result of $\mathcal{A}$ for each bucket $z\in\varphi(\textsc{Ball}(x,r))$. The analysis is straightforward and follows easily from Definition 5.2. Previously, an MPC version of Lemma 5.4 was introduced in [DBLP:conf/icalp/CzumajGJK024], but their algorithm is not efficient due to the lack of efficient consistent hashing which we devise (Lemma 5.3).